OpenStudy (anonymous):

What is the x-intercept of the line with the equation 8x-1/3y=16 PLZ HELP

OpenStudy (anonymous):

@PoyoPoyo @jane11509 @Chris01 ?? helpp?

OpenStudy (anonymous):

Me am no math magician. Sorry Kayla, its been like 3 years since I've done this math :P

OpenStudy (anonymous):

lol ok thanks anywats :P

OpenStudy (anonymous):

anyways*

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

Is it \[\frac{ 8x-1 }{ 3y }=16 ~~~ or ~~~ 8x-\frac{ 1 }{ ?3y}=16\]

OpenStudy (anonymous):

i have noo clue, all i can do is enter 3 numbers/letters in the blank D:

OpenStudy (anonymous):

Okay... can u send me a pic of your question?

OpenStudy (anonymous):

i can try:) one second

OpenStudy (anonymous):

Ya! :)

OpenStudy (anonymous):

it wouldnt let me :( but heres what it looked like

OpenStudy (anonymous):

Okay I can understand it now!

OpenStudy (anonymous):

Haha Okay!

OpenStudy (anonymous):

When you solve the equation, 8x = 16 + 1/3y => x = 2 + 1/24y

OpenStudy (anonymous):

Oh okay I think i got it now, Thanks!

OpenStudy (anonymous):

x intercept is a point on the graph where y = 0

OpenStudy (anonymous):

You're welcome! :)

OpenStudy (anonymous):

:)

OpenStudy (testing123guest):

@AlexandervonHumboldt2 so whats the answer here

OpenStudy (testing123guest):

@Michele_Laino do u know what the answer here is

OpenStudy (testing123guest):

@triciaal plz say u know the answer here

OpenStudy (michele_laino):

we get the y-intercept, if we replace \(y=0\) into the starting equation

OpenStudy (testing123guest):

so it would be ( y,0)

OpenStudy (triciaal):

the x-intercept is the value of x when y = 0

OpenStudy (testing123guest):

so the answer is ( 2 , 0 )

OpenStudy (triciaal):

@Testing123guest correct

OpenStudy (testing123guest):

cool thx

OpenStudy (triciaal):

welcome